# Displacement-Squeeze receiver for BPSK displaced squeezed vacuum states surpassing the coherent-states Helstrom bound under imperfect conditions

### v1

@@ III.2 Error probability under ideal conditions
For S-BPSK, the Helstrom bound in Eq. (9) and the DSR error probability in Eq. (20) satisfy
$\frac{P_{\text{err}}^{\text{DSR}}}{P_{\text{HB}}^{\text{DSS}}}=\frac{\exp\left\{-4N\left(N+1\right)\right\}}{1-\sqrt{1-\exp\left\{-4N(N+1)\right\}}}$ (21)
@@
In the low-energy limit $N\to 0$, both $P_{\text{err}}^{\text{DSR}}$ and $P_{\text{HB}}^{\text{DSS}}$ approach $0.5$ and the ratio tends to $1$, corresponding to random guessing. In the high-energy regime $N\gg 1$, the ratio approaches $2$ (i.e., $3$ dB), yielding
$P_{\text{err}}^{\text{DSR}}\simeq 2P_{\text{HB}}^{\text{DSS}}$ (22)
@@
Although the DSR remains suboptimal with respect to the Helstrom bound of S-BPSK and satisfies $P_{\text{err}}^{\text{DSR}}\in\left[P_{\text{HB}}^{\text{DSS}},2P_{\text{HB}}^{\text{DSS}}\right]$ for all $N$, it still provides a substantial advantage over C-BPSK channels. In particular, for the same mean photon number $N$, the proposed DSR can achieve an error probability below the Helstrom bound $P_{\text{HB}}^{\text{CS}}$ of a C-BPSK channel; for example, around $N\approx 0.4$ in Fig. 4 the DSR performance lies between the DSS Helstrom bound and the coherent state Helstrom bound.
@@
Figure 4: Performance of the DSR with an SPD under ideal conditions. (a) Error probability of the DSR. (b) Ratio (in dB) of the benchmarks $P_{\text{HB}}^{\text{CS}}$, $P_{\text{SQL}}^{\text{CS}}$, $P_{\text{HB}}^{\text{DSS}}$, and $P_{\text{SQL}}^{\text{DSS}}$ relative to $P_{\text{err}}^{\text{DSR}}$.
Figure 4(a) shows the DSR error probability as a function of the mean photon number. The DSR surpasses the C-BPSK SQL, the S-BPSK SQL, and the C-BPSK Helstrom bound at $N\approx 0.21$, $0.30$, and $0.40$, respectively. As $N$ increases, the performance advantage over these benchmarks grows. At $N=1.0$, the DSR achieves gains of $21.3$ dB, $11.4$ dB, and $14.4$ dB relative to the C-BPSK SQL, the S-BPSK SQL, and the C-BPSK Helstrom bound, respectively, while remaining $3$ dB above the S-BPSK Helstrom bound, as summarized in Fig. 4(b). These results demonstrate that the DSR provides strong performance over a broad photon-number regime.
@@
IV The performance of DSR under realistic conditions
IV.1 The quantum efficiency $\eta$ and the dark count rate $\nu$ of PNR(M)
The realistic factors of PNR detectors include quantum efficiency $\eta$ and dark count rate $\nu$. Quantum efficiency essentially corresponds to photon loss during detection, while dark count rate fundamentally arises from random triggering events caused by environmental noise [12, 7]. Under ideal conditions, quantum efficiency $\eta=1$ and dark count rate $\nu=0$. If we simultaneously consider imperfect quantum efficiency $\eta<1$ and dark count rate $\nu>0$, the POVM operators for PNR(M) (Eq. (14)) are rewritten as [2]:
$\Pi_{n}=\mathrm{e}^{-\nu}\sum_{l=0}^{n}\sum_{k=n-l}^{\infty}\frac{\nu^{l}}{l!}\mathrm{C}_{n-l}^{k}\cdot\eta^{n-l}(1-\eta)^{k-(n-l)}\left|k\right\rangle\left\langle k\right|,\ n=0,...,M$ (23)
At this point, the probability of PNR(M) detecting $n$ photons (Eq. (15)) is rewritten as:
@@ IV.1 The quantum efficiency $\eta$ and the dark count rate $\nu$ of PNR(M)
$q_{n}^{M}\left(\mu,\eta,\nu\right)=\left\{\begin{aligned} &\mathrm{e}^{-(\eta\mu+\nu)}\cdot\dfrac{\left(\eta\mu+\nu\right)^{n}}{n!},\quad n\leq M-1\\ &1-\mathrm{e}^{-(\eta\mu+\nu)}\cdot\sum_{j=0}^{M-1}\dfrac{\left(\eta\mu+\nu\right)^{j}}{j!},\quad n=M\end{aligned}\right$. (24)
By substituting into Eq. (17), the optimal threshold $n_{th}^{*}$ is obtained:
$n_{\text{th}}^{*}=\min\left\{\left\lceil\frac{\eta|\gamma|^{2}}{\ln\left(\eta|\gamma|^{2}+\nu\right)-\ln\nu}\right\rceil,M\right\}=\min\left\{\left\lceil\frac{\eta\cdot 4N(N+1)}{\ln\left(\eta\cdot 4N(N+1)+\nu\right)-\ln\nu}\right\rceil,M\right\}$ (25)
If only consider quantum efficiency ($\eta<1,\ \nu=0$), the optimal threshold $n_{th}^{*}$ remains 1 ($\frac{\eta|\gamma|^{2}}{\ln\left(\eta|\gamma|^{2}+\nu\right)-\ln\nu}\to 0^{+},\ \left\lceil\frac{\eta|\gamma|^{2}}{\ln\left(\eta|\gamma|^{2}+\nu\right)-\ln\nu}\right\rceil\to 1$). Then, the error probability expression (Eq. (20)) is rewritten as:
@@
$P_{\text{err}}^{\text{DSR}}(N,\eta)$ $=1-\frac{1}{2}\left[q_{0}^{M}(0,\eta)+1-q_{0}^{M}\left(|\gamma|^{2},\eta\right)\right]$ (26)
$=\frac{1}{2}\exp\left\{-4|\alpha|^{2}\cdot\mathrm{e}^{2r}\cdot\eta\right\}=\frac{1}{2}\exp\left\{-4N\left(N+1\right)\cdot\eta\right\}$
We can see that under non-ideal quantum efficiency condition, the resolution M of the PNR(M) in the DSR does not affect the error probability. Fig. 5(a) illustrates the variation of $P_{\text{err}}^{\text{DSR}}(N,\eta)$ (Eq. (26)) with respect to the average signal photon number N for $\eta$ = 1.0, 0.9, and 0.8. As shown, the three curves surpass the SQL of S-BPSK at N = 0.30, 0.39, and 0.56, respectively. Moreover, as N increases, the performance advantage of $P_{\text{err}}^{\text{DSR}}(N,\eta)$ over the SQL becomes more pronounced. At N = 1.0, the ratios of the three curves to the SQL are -4.5 dB, -7.97 dB, -11.44 dB, respectively.
Likewise, if only consider dark counts ($\eta=1,\ \nu>0$), the optimal threshold $n_{th}^{*}$ reads:
$n_{\text{th}}^{*}=\min\left\{\left\lceil\frac{|\gamma|^{2}}{\ln\left(|\gamma|^{2}+\nu\right)-\ln\nu}\right\rceil,M\right\}=\min\left\{\left\lceil\frac{4N(N+1)}{\ln\left(4N(N+1)+\nu\right)-\ln\nu}\right\rceil,M\right\}$ (27)
As shown, under non-ideal dark count conditions, $n_{\text{th}}^{*}$ is not always unity and varies with the average photon number (see Fig. 5(c)). In other words, in non-ideal dark count condition, the error probability of the DSR is correlated with the resolution of the PNR.
Figure 5: Performance of the DSR with an non-ideal PNR(M) which quantum efficiency $\eta<1$ and dark count $\nu>0$. (a) and (b) Error probability of the DSR. (c) Optimal threshold $n_{\text{th}}^{*}$. (d) Ratio (in dB) of the benchmark $P_{\text{SQL}}^{\text{DSS}}$ relative to $P_{\text{err}}^{\text{DSR}}$.
Fig. 5(b) shows the error probability as a function of the average signal photon number under two scenarios: considering only non-ideal dark counts, and considering both non-ideal detection efficiency and dark counts. In the presence of dark counts, the error probability exhibits an oscillatory pattern, which results from the continuous increase of the optimal threshold. This optimal threshold is limited by the resolution of the PNR detector; once it reaches the maximum resolvable photon number, the error probability no longer decreases with increasing average photon number but instead saturates. Furthermore, we observe that for curves with the same dark count rate but different detection efficiencies, the saturated error probabilities coincide. In contrast, for curves with the same detection efficiency but different dark count, a lower dark count leads to a lower saturated error probability.
We therefore conclude that the saturated error probability is determined primarily by the dark count, and its effect cannot be mitigated by increasing the signal strength. On the other hand, the performance degradation caused by non-ideal detection efficiency can be compensated by using a stronger signal. For practical PNR detectors, high detection efficiency and low dark count rate generally exhibit a trade-off relationship; improving the detection efficiency typically increases the dark count [30, 19]. Hence, in actual experiments, it can be beneficial to moderately reduce the detection efficiency in order to obtain a lower dark count, while compensating for the resultant loss in sensitivity by increasing the signal power.
In addition, a high dark count forces the optimal threshold to increase frequently until it reaches the resolution limit of the PNR detector, thereby driving the DSR scheme into saturation (see Fig. 5(c)). Consequently, a configuration with lower detection efficiency and lower dark count can alleviate the requirement for a PNR detector with very high resolution. Fig. 5(d) plots the ratio (in dB) of the error probability to the SQL of S-BPSK. It can be seen that even with both non-ideal detection efficiency and dark counts, the DSR scheme still surpasses the SQL of S-BPSK. For example, when $\eta=0.9$ and $\nu=10^{-2}$, the error probability of DSR with PNR(M) is lower than the SQL by -0.96dB (M=1), -3.07dB (M=2), -6.04dB (M=3), -9.58dB (M=4), and -13.52dB (M=5).
IV.2 Phase Diffuse Noise in communication channel
@@
Phase diffusion noise is a common communication channel noise that increases the overlap between states (S-BPSK), making them more difficult to distinguish [20, 6]. A schematic of a phase diffusion noise channel is shown in Fig. 6.
@@
Figure 6: Schematic diagram of phase diffusion noise channel. The signal state $\rho_{i}$ (pure state) is transformed into the state $\rho_{i}^{pd}$ (mixed state) in the phase diffuse communication channel.
@@
After DSSs are contaminated by phase diffusion noise in the channel, they evolve from pure states to mixed states and can only be expressed using density operators. Let the uncontaminated DSS be denoted as $\rho_{i}=\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|$. The expression for the contaminated quantum state $\rho_{i}^{\text{pd}}$ is then:
@@
$\rho_{i}^{\text{pd}}$ $=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g_{\sigma}(\phi)\cdot\rho_{i,\phi}\right\}$ (28)
$=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g_{\sigma}(\phi)\cdot R_{\phi}\rho_{i}R_{\phi}^{\dagger}\right\}$
@@
$\rho_{i,\phi}=\left|\psi_{i,\phi}\right\rangle\left\langle\psi_{i,\phi}\right|$ (29)
$\left|\psi_{i,\phi}\right\rangle=R_{\phi}\left|\psi_{i}\right\rangle=D\left((-1)^{i+1}\alpha\mathrm{e}^{-j\phi}\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$
@@
Figure 7: Phase-space representations of S-BPSK signal (mean photon number $N=1.0$ and phase diffusion strength $\sigma=0.5$). (a) DSSs $\rho_{i}$ without phase diffusion noise; (b) states $\rho_{i}^{\text{pd}}$ with phase diffusion noise; (c) states $\zeta_{i}$ after DSR processing.
For the DSR receiver, since the input state $\rho_{i}^{\text{pd}}$ is no longer Gaussian, the quantum state $\zeta_{i}$ after displacement-squeezing processing and prior to PNR detection also ceases to be Gaussian.
$\zeta_{i}$ $=U\rho_{i}^{\text{pd}}U^{\dagger}$ (30)
$=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g(\phi)\cdot UR_{\phi}\rho_{i}R_{\phi}^{\dagger}U^{\dagger}\right\}$
@@
$\left|\psi_{0,\phi}^{\prime}\right\rangle=U\cdot\left|\psi_{0,\phi}\right\rangle=S\left(-r\right)D\left(-\alpha\mathrm{e}^{-j\phi}+\alpha\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$ (31)
$\left|\psi_{1,\phi}^{\prime}\right\rangle=U\cdot\left|\psi_{1,\phi}\right\rangle=S\left(-r\right)D\left(+\alpha\mathrm{e}^{-j\phi}+\alpha\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$
@@
$q_{n}^{M}$ $\left(\zeta_{i}\right)=\mathbf{Tr}\left\{\Pi_{n}\cdot\zeta_{i}\right\}$ (32)
@@
$=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot\mathbf{Tr}\left\{\Pi_{n}\cdot\left|\psi_{i,\phi}^{\prime}\right\rangle\left\langle\psi_{i,\phi}^{\prime}\right|\right\}$
By substituting into Eq. (17) and Eq. (19), we obtain the optimal threshold and the error probability of the DSR under phase diffusion noise, respectively.
@@
Figure 8: Simulation results of DSR under phase diffusion noise ($\sigma=0.1$): (a) Error probability, (b) Ratio (in dB) with a benchmark of $P_{\text{SQL}}^{\text{DSS}}$ (Eq. (45)), (c) Optimal threshold $n_{th}^{*}$ and (d) Fock-basis population of measured states $\zeta_{i}=U\rho_{i}^{\text{pd}}U^{\dagger}$ by PNR(M), which $N=1.0$. Note: the SQL and HB of S-BPSK under phase diffusion condition is rewritten as Eq. (45) and Eq. (47).
Fig. 8(c) shows the variation of the optimal threshold $n_{\text{th}}^{*}$ for DSR with the mean photon number under a phase diffusion of $\sigma=0.1$. In the absence of resolution limits from the PNR detector, $n_{\text{th}}^{*}$ increases in steps of 2, which differs from the behaviors previously analyzed for dark counts and thermal noise. This is attributed to the fact that, in the presence of phase diffusion noise, DSR cannot transform $\rho_{0}^{\text{pd}}$ into the vacuum state or a coherent state with a very small mean photon number (see Fig. 7(c)). After DSR processing, the Fock-basis population of the resulting state $\zeta_{i}$ shows a feature reminiscent of squeezed states: the populations of certain even photon numbers become higher than those of adjacent odd photon numbers, as illustrated in Fig. 8(d).

### v2

@@ II Channel model and performance benchmarks
For S-BPSK, the Helstrom bound in Eq. (9) and the IS-Kennedy error probability in Eq. (22) satisfy the relation:
$\frac{P_{\text{err}}^{\text{ISK},\text{ideal}}}{P_{\text{HB}}^{\text{S-BPSK}}}=\frac{\exp\left\{-4N\left(N+1\right)\right\}}{1-\sqrt{1-\exp\left\{-4N(N+1)\right\}}}.$ (24)
@@
In the low-energy limit $N\to 0$, both $P_{\text{err}}^{\text{ISK}}$ and $P_{\text{HB}}^{\text{S-BPSK}}$ approach $0.5$, and their ratio approaches $1$, corresponding to random guessing. In the high-energy regime $N\gg 1$, the ratio approaches $2$ (i.e., $3$ dB), yielding:
$P_{\text{err}}^{\text{ISK},\text{ideal}}\simeq 2P_{\text{HB}}^{\text{S-BPSK}}.$ (25)
@@
Although the IS-Kennedy remains suboptimal with respect to the S-BPSK Helstrom bound, satisfying $P_{\text{err}}^{\text{ISK}}\in\left[P_{\text{HB}}^{\text{S-BPSK}},2P_{\text{HB}}^{\text{S-BPSK}}\right]$ for all $N$, it provides a substantial advantage over C-BPSK channels. Specifically, for the same mean photon number $N$, the proposed IS-Kennedy achieves an error probability strictly lower than the Helstrom bound $P_{\text{HB}}^{\text{C-BPSK}}$ of a C-BPSK channel. For instance, around $N\approx 0.4$ in Fig. 4, the IS-Kennedy performance lies between the S-BPSK Helstrom bound and the C-BPSK Helstrom bound.
@@
Figure 4: Performance of the IS-Kennedy under ideal conditions. (a) Error probability of the IS-Kennedy. (b) Ratio (in dB) of the benchmarks $P_{\text{HB}}^{\text{C-BPSK}}$, $P_{\text{SQL}}^{\text{C-BPSK}}$, $P_{\text{HB}}^{\text{S-BPSK}}$, $P_{\text{SQL}}^{\text{S-BPSK}}$, and $P_{\text{err}}^{\text{K},\text{ideal}}$ relative to $P_{\text{err}}^{\text{ISK},\text{ideal}}$.
Figure 4(a) shows the IS-Kennedy error probability as a function of the mean photon number $N$ under ideal conditions, alongside the Kennedy receiver and benchmarks for C-BPSK and S-BPSK discrimination. The IS-Kennedy surpasses the C-BPSK SQL, the S-BPSK SQL, and the C-BPSK Helstrom bound at approximately $N\approx 0.21$, $0.30$, and $0.40$, respectively. The performance advantage over these benchmarks increases with $N$, consistent with the fact that $P_{\text{err}}^{\text{ISK}}=\frac{1}{2}\exp[-4N(N+1)]$ features an approximately quadratic exponent $N(N+1)$, leading to a steeper decay than the linear scaling of coherent-state benchmarks. The dB ratios summarized in Fig. 4(b) quantify these advantages. At $N=1.0$, the IS-Kennedy achieves gains of $21.3$ dB, $11.4$ dB, and $14.4$ dB relative to the C-BPSK SQL, the S-BPSK SQL, and the C-BPSK Helstrom bound, respectively. Additionally, the IS-Kennedy significantly outperforms the conventional Kennedy receiver, exhibiting a $17.4$ dB gain at $N=1.0$. Meanwhile, it remains $3$ dB above the S-BPSK Helstrom bound, staying within a constant factor of the ultimate quantum limit. These results demonstrate that the IS-Kennedy delivers robust performance over a broad photon-number regime under ideal conditions.
@@
IV The performance of IS-Kennedy under imperfect conditions
IV.1 Imperfect PNR detector with detection efficiency, dark count, and finite resolution
The previous section presented the closed-form error probability of the IS-Kennedy receiver under ideal detection conditions, pointing out that with perfect nulling and matched squeezing, the optimal decision rule degenerates to a simple on–off strategy. However, in practical implementations, PNR detectors possess finite resolution (maximum resolvable photon number, M) and suffer from non-ideal factors such as limited detection efficiency $\eta$ and dark counts $\nu$ (background counts). In this section, while maintaining the assumption of an ideal IS-Kennedy processing module (i.e., retaining $|\zeta_{0}=\ket{0}\bra{0}$ and $|\zeta_{1}=\ket{2\gamma}\bra{2\gamma}$), we analyze the impact of imperfect PNR detection on the error probability and derive the expression for the optimal threshold (MAP-equivalent) decision and the corresponding error probability.
Considering an imperfect detection efficiency $\eta\in(0,1]$ and a mean dark count rate $\nu\geq 0$, the POVM operators are described by [3, 19]:
$|.\Pi_{n}^{\left(\eta,\nu\right)}=\mathrm{e}^{-\nu}\sum_{l=0}^{n}\sum_{k=n-l}^{\infty}\frac{\nu^{l}}{l!}\binom{k}{n-l}\eta^{n-l}(1-\eta)^{k-(n-l)}\ket{k}\bra{k}$. (26)
Taking the PNR detector’s resolution M into account (denoted as PNR(M)), the POVM of a PNR(M) detector is
@@ IV.1 Imperfect PNR detector with detection efficiency, dark count, and finite resolution
$\Pi_{n}^{\left(\eta,\nu,\text{M}\right)}=\left\{\begin{aligned} &\Pi_{n}^{\left(\eta,\nu\right)},\quad n=0,1,\cdots,\text{M}-1,\\ &\sum_{m=M}^{\infty}\Pi_{m}^{\left(\eta,\nu\right)},\quad n=\text{M}.\end{aligned}\right$. (27)
Consequently, the probability of detecting $n$ photons (Eq. (14)) is reformulated as:
$P(n|i)=\text{Tr}\left\{\zeta_{i}\Pi_{n}^{\left(\eta,\nu,\text{M}\right)}\right\}=\left\{\begin{aligned} &\text{Poiss}\left(n;\eta\mu_{i}+\nu\right),\quad n<\text{M}-1,\\ &\sum_{m=M}^{\infty}\text{Poiss}\left(m;\eta\mu_{i}+\nu\right),\quad n=\text{M},\end{aligned}\right$. (28)
@@
where $\mu_{0}=0$ and $\mu_{1}=\left|2\gamma\right|^{2}=4N_{\text{eff}}$. The log-likelihood ratio (Eq. (17)) is updated to:
$L(n)=-\eta\mu_{0}+\eta\mu_{1}+n\ln\frac{\eta\mu_{0}+\nu}{\eta\mu_{1}+\nu},$ (29)
which remains a monotonic function of $n$. Therefore, the MAP criterion is equivalent to a threshold detection scheme. Therefore, the optimal threshold (Eq. (19)) involves a truncation at M:
$n_{\text{th}}^{*}=\min\left\{\left\lceil\frac{\eta\mu_{1}-\eta\mu_{0}}{\ln\left(\eta\mu_{1}+\nu\right)-\ln\left(\eta\mu_{0}+\nu\right)}\right\rceil,\text{M}\right\}=\min\left\{\left\lceil\frac{4\eta N_{\text{eff}}}{\ln\left(4\eta N_{\text{eff}}+\nu\right)-\ln\nu}\right\rceil,\text{M}\right\}.$ (30)
From the above equation, the adaptive behavior of the IS-Kennedy receiver against different non-ideal factors can be observed:
  • Presence of dark counts ($\nu>0$): The denominator decreases, resulting in an optimal threshold $n_{\text{th}}^{*}>1$. This implies that the PNR detector can filter out background noise by raising the decision threshold, demonstrating its robustness advantage over a standard SPD.
  • Absence of dark counts ($\nu=0$): The denominator approaches infinity, leading to $n_{\text{th}}^{*}\to 1$. In this limit, the optimal strategy reverts to the standard on–off decision. This indicates that in a pure loss channel ($\eta<1,\nu=0$), the PNR detector offers no performance gain over an SPD, as simple presence/absence detection is optimal.
Figure 5: Performance analysis of the IS-Kennedy receiver under imperfect detection conditions. (a) Error probability of IS-Kennedy with SPD for varying detection efficiencies $\eta\in\left\{1.0,0.8,0.4\right\}$ and extremely small dark counts ($\nu=10^{-9}$). (b) Error probability of IS-Kennedy with PNR(2) under different dark count rates $\nu$ (with $\eta=1.0$). The “step-like” behavior in the error curves arises from the discrete nature of the optimal decision threshold. (c) The optimal detection threshold $n_{\text{th}}^{*}$ as a function of signal energy for different dark count rates, calculated via Eq. (30). The jumps in threshold correspond exactly to the inflection points observed in panels (b) and (d). (d) The performance ratio (in dB) of the IS-Kennedy with PNR(10) relative to the S-BPSK SQL. The oscillatory behavior reflects the discrete adjustment of $n_{\text{th}}^{*}$ to balance signal detection against dark count noise.
The parameter ranges considered in Fig. 5 are chosen to represent experimentally relevant classes of photon-number-resolving detectors. Near-unity efficiencies with negligible dark counts model state-of-the-art superconducting PNR detectors, such as segmented superconducting nanowire devices, for which system detection efficiencies approaching 98% with low dark-count rates have recently been reported in Ref. [13]. By contrast, a markedly lower-efficiency regime, around $\eta\approx 0.4$, is representative of silicon photomultiplier (SiPM) or multi-pixel photon-counter platforms, which have been successfully characterized and exploited in quantum-optics experiments despite their lower nominal efficiency and stronger spurious effects such as dark counts and optical cross-talk [9, 8]. To address this experimentally relevant drastic scenario, we include an additional low-efficiency curve ($\eta=0.4$) in Fig. 5(a).
As shown in Fig. 5(b) and (d), the error probability and the gain (in dB) relative to the S-BPSK SQL exhibit a characteristic oscillatory or “step-like” behavior when dark counts are significant. This phenomenon is intrinsic to the discreteness of PNR detection. As the signal energy $N$ increases, the optimal likelihood ratio varies continuously, whereas the physical decision threshold $n_{\text{th}}^{*}$ must remain an integer. Consequently, $n_{\text{th}}^{*}$ stays constant over intervals of $N$ (see Fig. 5(c)) before discretely jumping to the next integer level. These jumps allow the receiver to adaptively filter out higher dark-count rates, maintaining robust performance even when $\nu$ is large (e.g., $\nu=10^{-2}$).
Notably, Fig. 5(a) reveals that while the detection efficiency $\eta$ impacts the rate of error decay (the slope), the saturation error floor at high energies is dictated solely by the dark count rate $\nu$. In the high-energy regime ($N\gg 1$), the miss detection probability (Eq. (20)) rapidly approaches zero. Consequently, the error probability (saturation floor) is dominated by the false alarm probability, with the optimal threshold saturating at $n_{\text{th}}^{*}=\text{M}$:
$P_{\text{sat}}\approx\frac{1}{2}P_{\text{FA}}=\frac{1}{2}\mathrm{e}^{-\nu}\sum_{n=\text{M}}^{+\infty}\frac{\nu^{n}}{n!}\xrightarrow{\text{Taylor series}}\frac{1}{2}\left(1-\nu+\cdots\right)\sum_{n=\text{M}}^{+\infty}\frac{\nu^{n}}{n!}.$ (31)
Given that the dark count rate of practical PNR detectors is typically much less than unity ($\nu\ll 1$[13], the dominant contribution to the summation comes from the first term, $n=\text{M}$. Thus, the saturation error probability can be approximated as:
$P_{\text{sat}}\approx\frac{1}{2}\frac{\nu^{\text{M}}}{\text{M}!}.$ (32)
This expression indicates that the saturation error probability of the IS-Kennedy receiver is independent of detection efficiency $\eta$ and mean photon number $N$, depending instead on the PNR detector’s resolution M and dark count rate $\nu$. Therefore, the existence of any non-zero dark count, however small, inevitably leads to an error probability saturation. Furthermore, in conjunction with Eq. (23), since the advantage of the IS-Kennedy equipped with an SPD lies in converting squeezing resources into effective energy gain rather than suppressing noise floor, its saturation error probability is identical to that of the conventional Kennedy receiver (see Fig. 5(a)).
IV.2 Phase Diffusion Noise in communication channel
@@
Phase diffusion noise is a common communication channel noise that increases the overlap between states, making them more difficult to distinguish [26, 10]. A schematic of a phase diffusion noise channel is shown in Fig. 6.
@@
Figure 6: Schematic diagram of phase diffusion noise channel. The signal state $\rho_{i}$ (pure state) is transformed into the state $\rho_{i}^{pd}$ (mixed state) in the phase diffusion communication channel.
Figure 7: Phase-space representations of S-BPSK signal (mean photon number $N=1.0$ and phase diffusion strength $\sigma=0.5$). (a) S-BPSK $\rho_{i}$ without phase diffusion noise; (b) states $\rho_{i}^{\text{pd}}$ with phase diffusion noise; (c) states $\zeta_{i}$ after IS-Kennedy processing.
@@
After S-BPSK signals are contaminated by phase diffusion noise in the channel, they evolve from pure states to mixed states and can only be expressed using density operators. Let the uncontaminated signal be denoted as $\rho_{i}=\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|$. The expression for the contaminated quantum state $\rho_{i}^{\text{pd}}$ is then:
@@
$\rho_{i}^{\text{pd}}=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g_{\sigma}(\phi)\cdot\rho_{i,\phi}\right\}=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g_{\sigma}(\phi)\cdot R_{\phi}\rho_{i}R_{\phi}^{\dagger}\right\}$ (33)
@@
$\rho_{i,\phi}\sim\left|\psi_{i,\phi}\right\rangle=R_{\phi}\left|\psi_{i}\right\rangle=D\left((-1)^{i+1}\alpha\mathrm{e}^{-j\phi}\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$ (34)
@@
For the IS-Kennedy receiver, since the input state $\rho_{i}^{\text{pd}}$ is no longer Gaussian, the quantum state $\zeta_{i}$ after displacement-squeezing processing and prior to PNR detection also ceases to be Gaussian.
$\zeta_{i}$ $=U\rho_{i}^{\text{pd}}U^{\dagger}=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot\left\{g(\phi)\cdot UR_{\phi}\rho_{i}R_{\phi}^{\dagger}U^{\dagger}\right\}$ (35)
@@
$\left|\psi_{i,\phi}^{\prime}\right\rangle=U\cdot\left|\psi_{i,\phi}\right\rangle=S\left(-r\right)D\left((-1)^{i}\alpha\mathrm{e}^{-j\phi}+\alpha\right)S\left(r\mathrm{e}^{-j2\phi}\right)\left|0\right\rangle$ (36)
@@
$P(n|i)=\text{Tr}\left\{\Pi_{n}^{\text{ideal}}\cdot\zeta_{i}\right\}=\int_{-\infty}^{+\infty}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot\text{Tr}\left\{\Pi_{n}^{\text{ideal}}\cdot\left|\psi_{i,\phi}^{\prime}\right\rangle\left\langle\psi_{i,\phi}^{\prime}\right|\right\}$ (37)
@@
Figure 8: Simulation results of IS-Kennedy under phase diffusion noise ($\sigma=0.1$): (a) Error probability, (b) Ratio (in dB) with a benchmark of $P_{\text{SQL}}^{\text{S-BPSK}}$ (Eq. (65)), (c) Optimal threshold $n_{th}^{*}$ and (d) Fock-basis population of measured states $\zeta_{i}=U\rho_{i}^{\text{pd}}U^{\dagger}$ by PNR(M) for $N=1.5$. Note: the SQL and HB of S-BPSK under phase diffusion condition is rewritten as Eq. (65) and Eq. (67).
By substituting Eq. (37) into the MAP rule in Eq. (16), we numerically determine the optimal decision rule and the corresponding error probability under phase diffusion noise. We do not have a general closed-form proof that the likelihood ratio remains monotonic in $n$ for arbitrary $\sigma$; therefore, the threshold reported below should be understood as a numerically verified MAP-equivalent threshold over the explored parameter range.
Fig. 8(c) shows that the optimal threshold $n_{\text{th}}^{*}$ under phase diffusion ($\sigma=0.1$) exhibits a staircase-like dependence on the mean photon number $N$. In the absence of PNR resolution limits, $n_{\text{th}}^{*}$ increases predominantly in unit steps, while some even thresholds remain optimal only over very narrow intervals of $N$, leading to unequal step widths compared with the dark-count case in Sec. IV.A. This behavior originates from the non-Gaussianity induced by phase diffusion: after the IS-Kennedy processing, $\zeta_{i}$ cannot be perfectly nulled to (near-)vacuum and its photon-number distribution exhibits noticeable even-odd modulation (Fig. 8(d)). As $N$ varies, this modulation causes the likelihood function $L\left(n\right)$ to shift non-uniformly across adjacent photon numbers, so that certain even thresholds can be optimal only over short $N$ intervals. Consequently, $n_{\text{th}}^{*}$ still changes mainly by +1, but with uneven plateau lengths. We verified this by refining the sampling of $N$, which reveals short-lived plateaus at some even $n_{\text{th}}^{*}$ values that may be missed under coarser sampling.
Fig. 8(a) depicts the error probability of IS-Kennedy as a function of the mean signal photon number under a phase diffusion strength of $\sigma=0.1$. And Fig. 8(b) shows the ratio (in dB) relative to the S-BPSK SQL under $\sigma=0.01,0.1,0.2,1.0$. The results show that IS-Kennedy can surpass the SQL of the S-BPSK scheme within a certain photon-number range ($N\in\left(0.28,0.93\right)$ under $\sigma=0.1$), achieving a maximum advantage of 10.47 dB ($\sigma=0.01$), 3.32 dB ($\sigma=0.1$), 2.00 dB ($\sigma=0.2$) and 0.36 dB ($\sigma=1.0$) beyond the S-BPSK SQL.